# Machine learning for advancing low-temperature plasma modeling and simulation

### v1

@@ 2.2 Plasma chemistry
Low-temperature plasma models are based on a set of chemical reactions which describes volumetric interactions between all the species tracked in the model, such as electrons, ions, and neutrals.[61] An accurate description of such interactions is required in order to understand and optimize the selective generation of chemically active species and their properties, such as mean energy and velocity distribution. For plasma modeling applications, chemistry sets may include up to hundred species and several thousands of reactions.[62] Such chemical complexity can hinder the process of understanding and optimization of plasma applications. Furthermore, the plasma modeling community is sometimes facing scarce availability of fundamental data, such as cross sections and reaction rate coefficients, that are used as input in the models.[63] With significant advances of ML, new tools are available to the community to improve plasma chemistry models. There are three major roles of ML with respect to plasma chemistry developments, that are: (i) extraction of cross sections and reaction rate coefficients, (ii) derivation of reduced reaction mechanisms, and (iii) chemical engineering. In particular, the following questions are addressed in the present review:
How can electron-impact cross sections with atoms and molecules of interest for plasma applications be rapidly computed?
@@
The use of an ANN as an optimization technique for treating the inverse problem of obtaining electron collision cross section from electron transport coefficients has been proposed by Morgan.[64] In that pioneering work, it was assessed that neural networks are indeed useful to determine the elastic momentum transfer cross sections of electrons in real gases, but the accuracy level is highly dependent on the quality of input cross section data that are used for training of the model. Stokes and co-authors[65, 66] have improved the idea by Morgan and implemented deep neural network (DNN) using cross sections from the LXCat database[67] associated with electron transport coefficients found by numerical solutions of the electron Boltzmann equation or by experimental measurements. In their works, the authors demonstrated that their automatic solution using ANN had an accuracy comparable to that of a human expert in determining cross sections of electrons in He or in the biomolecule tetrahydrofuran (THF). In [66], they have also showed that the use of large amount of synthetic training data generated using the real cross sections available from LXCat gave good results for the prediction of elastic momentum transfer and ionization cross sections of He and Ar. However, the architecture of the ANN used in [65] and [66] had minor improvements over the architecture proposed by Morgan[64]. Jetly and Chaudhury [68] have implemented a ANN, specifically CNN and densely connected convolutional network (DenseNet) to derive electron impact cross sections from solution of the inverse electron swarm problem. The networks were trained using elastic momentum transfer, ionization, and excitation cross sections for different gases available in the LXCat database[67] and their corresponding electron transport coefficients, which were calculated using the BOLSIG+ solver.[69] For the first time, the performance and accuracy in cross sections determination from different network architectures was assessed and it was found that DenseNet predicts cross sections with significantly higher accuracy compared to other ANN models due to its ability to extract both long and short term features from the electron transport parameters. Figure 1 shows that the DenseNet model, as opposed to other CNN or ANN-based models, is able to predict specific cross sections features in energy, such as the resonance of the elastic momentum transfer cross sections of electrons in N.
@@
Direct determination of reaction rate coefficients and their temperature dependence is also important in plasma modeling. In fact, each reaction needs to be assigned with its own kinetic parameters and these are sometimes not available or subject of large uncertainties. Reiser and co-authors [70] have used ML to derive rate coefficients for radio-frequency atmospheric pressure plasmas operated with various gas mixtures of O, CH, and He. In their work, the authors used a genetic algorithm to derive rate coefficient values that bring the results of the integrated set of particle balance equations for different species in the plasma in the best possible agreement with the experimental data. Even if this approach could be prone to biases induced by user experience and be limited by the availability of experimental data, it has been demonstrated that results from the genetic algorithm open up the possibility of learning more about the significance and correlation of individual reactions. Hence, this method is potentially interesting for post-processing and evaluation of experimental data using simplified reaction schemes. Hanicinec and co-authors [71] have explored the use of ML to supply unknown reaction rates in plasma chemistries, thus allow complete chemistry sets to be generated without resorting to estimation by analogy or an educated guess. In their work, ML was used to provide rate coefficients of binary chemical reactions using three distinct optimized regression models: Support Vector Machine (SVM) regression model, the Random Forest regression model, and the Gradient-boosted Trees regression model. The regression models were trained on kinetic data for binary heavy-species collisions at or near room-temperature extracted from the QBD [72], KIDA [73], NFRI [74], and UfDA [75] databases. After removing duplicate reactions, the final data set consisted of 9470 reactions involving 1080 distinct species. As a sample use case, the ML results were used to augment the chemistry of a BCl/H gas mixture. In this example, reactions between the various BCl species and H were missing. Results of their work showed the importance of taking such reactions into consideration for which the ML algorithm likely gives better estimates than intuitive guesses, while being faster than measurements or ab initio calculations.[71]
@@
Graph theory has been applied to chemical kinetics for visualization, topological statistics calculation of a chemical network, and determination of pathways from a sequence of reactions. Sakai and co-authors[80] have been the first to apply it to complex plasma chemistry. In their work, they have constructed a network structure for methane and silane plasmas by representing the species as nodes and reactions as directed unweighted edges and they demonstrated that this network analysis allows one to estimate substantial effects of given species on other ones existing in the network. Then, the work was extended with betweenness and centrality indices to highlight not only the influentiality of each species on/to the others, but also global structures in chemical networks.[81] In this respect, these works have opened up possibilities for simpler prediction of chemical reactions that we would like to understand and/or control. A similar approach for graph construction has been adopted by Sun and co-authors,[82] with a different choice of undirected edges weighted based on steady-state reaction rates. The reduction method has been applied to CO chemical kinetics for a gliding arc plasma. Murakami and Sakai[83] have used a graph-based approach for extracting a low-dimensional reaction set based on centrality indices. In particular, their approach wa applied to HeO and Hehumid air plasma chemistry and they have shown that a reduced chemical reaction set for these systems can be achieved with the help of topological centrality (closeness and betweenness) and scale-freeness. In particular, in Figure 2, it is shown that a plasma-induced reaction network can be described with a scale-free network, where the number of degrees originating from a particular species follows a power-law distribution.
@@
The advantage of their approach is that it can be used as preprocessing to reduce complex plasma chemistry before multi-dimensional simulations. Mizui and co-authors[84] have extended the analysis in [83] to transient states by constructing temporal graphs. Their classification method was used to deduce information from silane and methane plasma chemistries. In particular, quantitative insights into chemical mechanisms have been derived with axes of multi-centrality indices. Holmes and co-authors[85] used direct graph representation of plasma mechanisms in which both species and reactions are identified as nodes, with the edges weighted based on the correspondent reaction rate coefficients. Furthermore, they have used a connectivity matrix to get qualitative information on fast and slow reaction pathways. With this approach, they have showed the new approach can produce a clearer image for the interpretation of complex chemistry. Venturi and co-authors[86] have presented a novel method for reducing plasma chemistry mechanisms by taking into account the uncertainties affecting the reaction parameters. As a difference with respect to previous works, their approach relies on an ensemble of weighted directed graphs having both species and reactions as nodes. Edges are weighted based on the reaction contribution to the particle balance equation. The methods were applied to a plasma mixture of 23 species including 15 vibrational states of the H molecule, three electronic states of the H atom, four ions, and electrons. As a result of their work, it has been shown that a graph ensemble significantly improves the methodology’s robustness while preserving the predictive accuracy. In general, graph theory related algorithm have been successfully applied to find patterns in large chemical data set. These algorithms could be used to identify and analyze patterns in chemical kinetic data prior running experiments or simulations [86].
@@
In many applications, the plasma community is interested in knowing “how we can” generate a specific concentration of species. In this respect, modeling and simulations are powerful tools to allow one to actively modulate plasmas and/or discharges for production of chemically reactive species. In this subsection, advances in ML for chemical engineering of plasma systems are explored. Zhu and co-authors[87] have developed a DL framework called DeePlasKin for non-equilibrium plasma systems that can be described by a global chemistry model. The model is based on a predictor-corrector approach detailed in subsection 2.4. For a given kinetic scheme and pre-defined temporal evolution of target species densities, it allows one to extract the temporal profile of the reduced electric field () and all other species densities. Pan and co-authors[88] have investigated a multi-layer feed-forward neural network for plasma kinetic modeling. The deep learning model has been applied to kinetic schemes for CHAr pulsed discharge and a NH plasma catalysis model for ammonia production. The magnitude of of the discharges and time were set as the input data of the DNN, whereas the target produce densities of each species calculated at every time from the kinetic model were set as output data. The results showed that the network can replace complex kinetic models with large reactions set, while significantly improving the computational efficiency. Wang and co-authors[49] have proposed an MLP with graphics processing unit (GPU) acceleration to describe the discharge characteristics and plasma chemistry of CO pulsed DBD discharges at atmospheric pressure. The MLP was trained using data obtained from fluid simulations. As a result of their work, the well trained model takes a few seconds to obtain various features of CO pulsed discharges, which significantly improves the computational efficiency with respect to the 30 hours CPU time that are required for the fluid model. Overall, these works have shown a great potential of ML for optimization and design of plasma sources in practical applications. In fact, the proposed models have been able not only to automatically extract plasma parameters that are difficult to measure in experiments, but also possess high computational efficiency, which provides a promising approach for multi-dimensional plasma modeling.
@@ 2.3 Plasma-surface interaction
Particle fluxes onto LTP bounding surfaces contain a variety of species (e.g., electrons, neutrals, ions, radicals) which trigger different kinds of PSIs upon impact, e.g., surface chemical reactions, sputtering. Each of which is more or less relevant for certain plasma applications, which are correspondingly engineered to lessen or strengthen one or the other mechanism. For modeling such processes and in particular the involved PSIs, a wide range of methodologies has been established, e.g., transport of ions in matter (TRIM)[89], molecular dynamics (MD). But each method is limited by a combination and trade off between accuracy and computational feasibility. Recent advances in both regards utilizing ML will be outlined in the following for processes that either are driven by (i) surface chemical reactions or (ii) physical sputtering.
@@
Plasmas introduce radicals and ions to ordinary gas-phase processing (e.g., chemical vapor deposition, atomic layer deposition, catalysis) and hereby enable a manifold of new reactions at the surface.
@@
This circumstance is employed for instance for the plasma enhanced atomic layer deposition (PEALD) of metal oxides. It consists of two self-limiting half cycles with purging cycles in between, out of which the second half cycle, the plasma cycle, is dedicated to the deposition of O atoms, molecules, and radicals. Ding et al.[90] have modeled the PEALD of HfO thin films with a kinetic Monte Carlo scheme. Reaction path ways and energy barriers were obtained by combining an intensive literature research with density functional theory (DFT) calculations. Such simulations were used to assemble a data set to train a Bayesian regularized artificial neural network (BRANN)[91] for the prediction of the half cycle process times, which are essential for industrial considerations, as a function of total precursor pressure and surface temperature. The BRANN was constructed by combing an MLP network architecture with a Bayesian regularization, which penalizes large weights to achieve a better generalization. In a followup work, [92] pair their kinetic Monte Carlo model with a computational fluid dynamics software to simulate the whole process. A RNN was then correspondingly trained to predict the evolution of the surface profile as well as gas-phase. Recurrent links establish internal states which allow for the description, prediction of the temporal dynamic behavior. The issue of vanishing as well as exploding gradients during the training of the RNN was addressed by including long short-term memory (LSTM, cf. subsection 3.1).
@@
The plasma enhanced atomic layer etching (PEALE) of copper has been suggested by Sheil et al.[93] to include a half cycle of plasma induced oxidation followed up by a half cycle of exposing the surface to etchant molecules (e.g., HCOOCH) which solely removes the just formed oxide layer. While the latter step is naturally self-limiting, the former is not. This finding has been seconded by Xia and Sautet,[94] who performed MD simulations for a corresponding plasma oxidation of copper. A high-dimensional neural network potential was trained by considering the ReaxFF potential[95, 96] and DFT calculations for the initial and on the fly data generation, respectively. The ML potential was demonstrated to overcome the limitations of classical as well as reactive interaction potentials imposed by their functional form, which are argued to not allow for accurate descriptions of very short range interatomic distances.
@@
However, for the particle fluxes from the plasma, not only radicals but molecules with high vibrational excitation are assumed to play an integral role for plasma catalysis. This is even although their contribution cannot be experimentally clarified due to an overlay with such effects as induced for instance by high electric fields. In contrast, simulations are well suited to provide experimentally unattainable insight. Wan et al.[97] have conducted DFT simulations to investigate the role of electric field-dipole interactions for the CH synthesis catalyzed with Ru surfaces. The resource intensive computations were accelerated by approximately five orders of magnitudes by generalizing the aggregated data with ANNs. Relevant atom configurations were processed with graph neural networks (GNNs),[98, 99] specifically graph convolutional networks,[100] whose network structure consists of nodes (i.e., atoms) and edges (i.e., configurational similarities). Multiple shared MLPs were used to combine the output with information on the electric field, adsorbate energy, and surface energy to predict all quantities of interest (i.e., adsorption energies, dipole moments, and polarizabilities of reactive intermediates). Kedalo et al.[101] have performed MD simulations to show that the activation of highly excited N molecules on Ru surfaces during NH synthesis can be approximated by the Fridman–Macheret model.[102] The interatomic interactions were described by a ML potential, utilizing the smooth overlap of the atomic positions (SOAP) descriptors,[103] to enable ab initio accuracy but with increased computational efficiency. Though Bal and Neyts[104] have demonstrated that the Fridman–Macheret model is insufficient to describe the vibrational nonequilibrium (i.e., lowering of the free energy barrier) during the catalytic dissociation of H and CH on Ni surfaces at low temperatures of the order of 750 K. The multitudes of potential reaction pathways in between two metastable states were described by simpler means by applying the harmonic linear discriminant analysis (supervised dimensionality reduction algorithm).[105] The states themselves were studied by conducting short MD simulations with enhanced sampling methods (i.e., variationally enhanced sampling[106], metadynamics[107, 108]) to introduce the vibrational nonequilibrium.
@@
Siron et al.[109] have proposed a high-throughput workflow for the study of amorphous material (i.e., a-C, a-Si, a-SiO, a-AlO) surface reactions relevant to dry plasma etching. Surface sites were characterized with the SOAP descriptor and clustered with a Bayesian Gaussian Mixture model to effectively reduce sites to be studied for reaction kinetics, that is etching.
@@
The bombardment of surfaces (e.g., targets, substrates) with energetic ions spawns collision cascades in the subsurface region, which eventually lead to persistent radiation damage and the emission of material, atoms, from the surface. The latter is referred to as sputtering and is of major importance for the sputter deposition of thin films as well as etching of high aspect ratio feature patterns (trenches).
@@
Hamedani et al.[110] have used a ML interaction potential to conduct a MD study on the radiation damage and sputtering of Si. The utilized ML potential, specifically the Gaussian approximation potential (GAP),[111] was originally trained to equilibrium properties and, hence, complemented with a DFT repulsive potential (i.e., DMol[112]) for short range interactions to resolve the evolution of the collision cascade accurately. The simulations provided sputtering yields that were in good agreement with experimental reference values. A greater number of isolated small defect clusters was observed than provided by MD simulations based on traditional interaction potentials (i.e., Stillinger-Weber[113], Tersoff III[114]). Moreover, a new defect structure was revealed, that is three split interstitials surrounding a central vacancy. This structure and the likelihood for its formation was further analyzed in a followup study.[115]
@@
Sputtering yields can also be predicted by utilizing the Yamamura model.[116] Phadke et al.[117] have showed that its model parameter Q, which relates to linear scaling, has a simple positive correlation with the sputter threshold energy. This result was obtained by applying a Bayesian Markov Chain Monte Carlo algorithm to TRIDYN[118] simulations of Ar and Ne ions bombarding 70 elemental targets. Eckstein and Preuss[119] applied Bayesian probability theory to improve the accuracy of the Yamamura model’s sputtering yield predictions by empirical means. Preuss et al.[120] conducted SDTrimSP[118, 121] simulations of D ions sputtering Fe targets to revise the model’s internal parameter (i.e., surface binding energy). This was achieved by to applying a reduced-order spectral expansion[122] (uncertainty quantification) based on the Bayesian framework.

### v2

@@ 2.2 Plasma chemistry
Low-temperature plasma models are based on a set of chemical reactions which describes volumetric interactions between all the species tracked in the model, such as electrons, ions, and neutrals.[61] An accurate description of such interactions is required in order to understand and optimize the selective generation of chemically active species and their properties, such as mean energy and velocity distribution. For plasma modeling applications, chemistry sets may include up to hundred species and several thousands of reactions.[62] Such chemical complexity can hinder the process of understanding and optimizing plasma applications. Furthermore, the plasma modeling community is sometimes facing scarce availability of fundamental data, such as cross sections and reaction rate coefficients, that are used as input in the models.[63] With significant advances of ML, new tools are available to the community to improve plasma chemistry models. There are three major roles of ML with respect to plasma chemistry developments, that are: (i) extraction of cross sections and reaction rate coefficients, (ii) derivation of reduced reaction mechanisms, and (iii) chemical engineering. In particular, the following questions are addressed in the present review:
How can electron-impact cross sections with atoms and molecules of interest be rapidly computed?
@@
The use of an ANN as an optimization technique for treating the inverse problem of obtaining electron collision cross section from electron transport coefficients has been proposed by Morgan.[64] In that pioneering work, it was assessed that neural networks are indeed useful to determine the elastic momentum transfer cross sections of electrons in real gases, but the accuracy level is highly dependent on the quality of input cross section data that are used for training of the model. Stokes and co-authors[65, 66] have improved the idea by Morgan and implemented deep neural network (DNN) using cross sections from the LXCat database[67] associated with electron transport coefficients found by numerical solutions of the electron Boltzmann equation or by experimental measurements. In their works, the authors demonstrated that their automatic solution using ANN had an accuracy comparable to that of a human expert in determining cross sections of electrons in He or in the biomolecule tetrahydrofuran (THF). The authors have also shown that the use of large amount of synthetic training data generated by using the real cross sections available from LXCat provide good results for the prediction of elastic momentum transfer and ionization cross sections of He and Ar.[66] However, the architecture of the ANN used in ref.[65] and ref.[66] had minor improvements over the architecture proposed by Morgan[64]. Jetly and Chaudhury [68] have implemented an ANN, specifically CNN and densely connected convolutional network (DenseNet)[37] to derive electron impact cross sections from a solution of the inverse electron swarm problem. The networks were trained using elastic momentum transfer, ionization, and excitation cross sections for different gases available in the LXCat database[67] and their corresponding electron transport coefficients, which were calculated using the BOLSIG+ solver.[69] For the first time, the performance and accuracy in cross sections determination from different network architectures was assessed and it was found that DenseNet predicts cross sections with significantly higher accuracy compared to other ANN models due to its ability to extract both long and short term features from the electron transport parameters. Figure 1 shows that the DenseNet model, as opposed to other ANN-based models (e.g., CNN), is able to predict specific cross sections features in energy, such as the resonance of the elastic momentum transfer cross sections of electrons in N.
@@
Fast prediction of collision cross sections based on simulated cross sections has been initially suggested by Zhong.[70] While the author used density functional theory (DFT) calculations to devise a data set for electron impact ionization cross sections for molecules, it has been argued that the approach would be similarly applicable to experimentally measured cross sections. Based on a support vector machine (SVM) model with radial basis function kernel, the prediction of accurate cross sections has been verified. Moreover, the capability to generalize from small to large molecules has been argued, without the need for computationally costly DFT calculations. Following a similar paradigm, Harris and Nepomuceno have recently developed a conceptually simple but effective ANN model to predict electron-molecule ionization cross sections from a database of measured cross sections.[71] In particular, by exploiting the chemical configuration (i.e., number of atoms in a molecule) as input to the network, the authors have evaluated the capability of a generalized prediction of cross sections for additional molecules.
Direct determination of reaction rate coefficients and their temperature dependence is also important in plasma modeling. In fact, each reaction needs to be assigned with its own kinetic parameters and these are sometimes not available or subject of large uncertainties. Reiser and co-authors [72] have used ML to derive rate coefficients for radio-frequency atmospheric pressure plasmas operated with various gas mixtures of O, CH, and He. In their work, the authors used a genetic algorithm to derive rate coefficient values that bring the results of the integrated set of particle balance equations for different species in the plasma into good agreement with the experimental data. Even if this approach could be prone to biases induced by user experience and be limited by the availability of experimental data, it has been demonstrated that results from the genetic algorithm open up the possibility of learning more about the significance and correlation of individual reactions. Hence, this method is potentially interesting for post-processing and evaluating experimental data using simplified reaction schemes. Hanicinec and co-authors [73] have explored the use of ML to supply unknown reaction rates in plasma chemistries, thus allow complete chemistry sets to be generated without resorting to estimations by analogy or an educated guess. In their work, ML was used to provide rate coefficients of binary chemical reactions using three distinct optimized regression models: An SVM model, a random forest model, and a gradient-boosted trees model. The models were trained on kinetic data for binary heavy-species collisions at or near room-temperature extracted from the QBD [74], KIDA [75], NFRI [76], and UfDA [77] databases. After removing duplicate reactions, the final data set consisted of 9470 reactions involving 1080 distinct species. As a sample use case, the ML results were used to augment the chemistry of a BCl/H gas mixture. In this example, reactions between the various BCl species and H were missing. Results of their work showed the importance of taking such reactions into consideration for which the ML algorithm likely gives better estimates than intuitive guesses, while being faster than measurements or ab initio calculations.[73]
@@
Graph theory has been applied to chemical kinetics for visualization, topological statistics calculation of a chemical network, and determination of pathways from a sequence of reactions. Sakai and co-authors[82] have been the first to apply it to complex plasma chemistry. In their work, they have constructed a network structure for methane and silane plasmas by representing the species as nodes and reactions as directed unweighted edges and they demonstrated that this network analysis allows one to estimate substantial effects of given species on other ones existing in the network. Then, the work was extended with betweenness and centrality indices to highlight not only the influentiality of each species on/to the others, but also global structures in chemical networks.[83] In this respect, these works have opened up possibilities for simpler prediction of chemical reactions that one would like to understand and/or control. A similar approach for graph construction has been adopted by Sun and co-authors,[84] with a different choice of undirected edges weighted based on steady-state reaction rates. The reduction method has been applied to CO chemical kinetics for a gliding arc plasma. Murakami and Sakai[85] have used a graph-based approach for extracting a low-dimensional reaction set based on centrality indices. In particular, their approach was applied to HeO and Hehumid air plasma chemistry and they have shown that a reduced chemical reaction set for these systems can be achieved with the help of topological centrality (closeness and betweenness) and scale-freeness. In particular, in Figure 2, it is shown that a plasma-induced reaction network can be described with a scale-free network, where the number of degrees originating from a particular species follows a power-law distribution.
@@
The advantage of their approach is that it can be used as preprocessing to reduce complex plasma chemistry before multi-dimensional simulations. Mizui and co-authors[86] have extended the analysis in ref.[85] to transient states by constructing temporal graphs. Their classification method was used to deduce information from silane and methane plasma chemistries. In particular, quantitative insights into chemical mechanisms have been derived with axes of multi-centrality indices. Holmes and co-authors[87] used a direct graph representation of plasma mechanisms in which both species and reactions are identified as nodes, with the edges weighted based on the corresponding reaction rate coefficients. Furthermore, they have used a connectivity matrix to get qualitative information on fast and slow reaction pathways. With this approach, they have shown that the new approach can produce a clearer image for the interpretation of complex chemistry. Venturi and co-authors[88] have presented a novel method for reducing plasma chemistry mechanisms by taking into account the uncertainties affecting the reaction parameters. As a difference with respect to previous works, their approach relies on an ensemble of weighted directed graphs having both species and reactions as nodes. Edges are weighted based on the reaction contribution to the particle balance equation. The methods were applied to a plasma mixture of 23 species including 15 vibrational states of the H molecule, three electronic states of the H atom, four ions, and electrons. As a result of their work, it has been shown that a graph ensemble significantly improves the methodology’s robustness while preserving the predictive accuracy. In general, graph theory related algorithms have been successfully applied to find patterns in large chemical data set. These algorithms could be used to identify and analyze patterns in chemical kinetic data prior running experiments or simulations [88].
@@
In many applications, the plasma community is interested in knowing “how we can” generate a specific concentration of species. In this respect, modeling and simulations are powerful tools to allow one to actively modulate plasmas and/or discharges for production of chemically reactive species. In this subsection, advances in ML for chemical engineering of plasma systems are explored. Zhu and co-authors[89] have developed a DL framework called DeePlasKin for non-equilibrium plasma systems that can be described by a global chemistry model. The model is based on a predictor-corrector approach detailed in subsection 2.4. For a given kinetic scheme and pre-defined temporal evolution of target species densities, it allows one to extract the temporal profile of the reduced electric field () and all other species densities. Pan and co-authors[90] have investigated a multi-layer feed-forward neural network for plasma kinetic modeling. The deep learning model has been applied to kinetic schemes for CHAr pulsed discharge and a NH plasma catalysis model for ammonia production. The magnitude of and time were chosen as the input data of the DNN, whereas the output was selected as the target densities of each species calculated at every instance of time by inquiring the kinetic model. The results showed that the network can replace complex kinetic models with large reactions set, while significantly improving the computational efficiency. Wang and co-authors[49] have proposed an MLP to describe the discharge characteristics and plasma chemistry of CO pulsed DBD discharges at atmospheric pressure. The MLP was trained using data obtained from fluid simulations. As a result of their work, the trained model takes a few seconds to predict various features of CO pulsed discharges, which significantly improves the computational efficiency with respect to the 30 hours CPU time that are required for the fluid model. Overall, these works have shown the great potential of ML for optimization and design of plasma sources in practical applications. In fact, the proposed models have been able not only to automatically extract plasma parameters that are difficult to measure in experiments, but also possess high computational efficiency, which provides a promising approach for multi-dimensional plasma modeling.
@@ 2.3 Plasma-surface interaction
Particle fluxes onto LTP bounding surfaces contain a variety of species (e.g., electrons, neutrals, ions, radicals) which trigger different kinds of PSIs upon impact, for instance surface chemical reactions or sputtering. Each of which is more or less relevant for certain plasma applications, which are correspondingly engineered to lessen or strengthen one or the other mechanism. For modeling such processes and in particular the involved PSIs, a wide range of methodologies has been established, such as transport of ions in matter (TRIM)[91] or molecular dynamics (MD). Each method is limited by a combination and trade off between accuracy and computational feasibility. Recent advances in both regards utilizing ML will be outlined in the following for processes that either are driven by (i) surface chemical reactions or (ii) physical sputtering.
@@
In addition to the species and surface chemical reactions inherent to ordinary gas-phase processing (e.g., chemical vapor deposition, atomic layer deposition, catalysis), plasmas introduce radicals, excited species, and ions and hereby enable a manifold of new reactions at the surface. This circumstance is employed for instance for the plasma enhanced atomic layer deposition (PEALD) of metal oxides. It consists of two self-limiting half cycles with purging cycles in between, out of which the second half cycle, the plasma cycle, is dedicated to the deposition of O atoms, molecules, and radicals. Ding et al.[92] have modeled PEALD of HfO thin films with a kinetic Monte Carlo scheme. Reaction path ways and energy barriers were obtained by combining an intensive literature research with DFT calculations. Such simulations were used to assemble a data set to train a Bayesian regularized artificial neural network (BRANN)[93]. The trained network was used for the prediction of the half cycle process times, which are essential for industrial considerations, as a function of total precursor pressure and surface temperature. The BRANN was constructed by combing an MLP network architecture with a Bayesian regularization, which penalizes large weights to achieve a better generalization. In a follow-up work, [94] the authors pair their kinetic Monte Carlo model with a computational fluid dynamics process simulation. An RNN was then correspondingly trained to predict the evolution of the surface profile as well as gas-phase kinetics. Recurrent links establish internal states which allow for the prediction of the temporal dynamics. The issue of vanishing as well as exploding gradients during the training of the RNN was addressed by including an LSTM (cf. subsection 3.1).
@@
Plasma enhanced atomic layer etching (PEALE) of copper has been suggested by Sheil et al.[95] to include a half cycle of plasma induced oxidation followed by a half cycle of exposing the surface to etchant molecules (e.g., HCOOCH), which solely removes the just formed oxide layer. While the latter step is naturally self-limiting, the former is not. This finding has been seconded by Xia and Sautet,[96] who performed MD simulations for a corresponding plasma oxidation of copper. A high-dimensional neural network potential was trained by considering the ReaxFF potential[97, 98] and DFT calculations for the initial and on-the-fly data generation, respectively. The ML potential was demonstrated to overcome the limitations of classical as well as reactive interaction potentials imposed by their functional form, which are argued to not allow for accurate descriptions of very short range interatomic distances.
@@
For the particle fluxes from the plasma, not only radicals but molecules with high vibrational excitation are assumed to play an integral role for plasma catalysis. This is assessed although their contribution cannot be experimentally clarified due to an overlay with effects, for instance as induced by high electric fields. In contrast, simulations are well suited to provide experimentally unattainable insight. Wan et al.[99] have conducted DFT simulations to investigate the role of electric field-dipole interactions for the CH synthesis catalyzed with Ru surfaces. The resource intensive computations were accelerated by approximately five orders of magnitudes by generalizing the aggregated data with ANNs. Relevant atom configurations were processed with graph neural networks (GNNs),[100, 101] specifically graph convolutional networks,[102] whose network structure consists of nodes (i.e., atoms) and edges (i.e., configurational similarities). Multiple shared MLPs were used to combine the output with information on the electric field, adsorbate energy, and surface energy to predict all quantities of interest (i.e., adsorption energies, dipole moments, and polarizabilities of reactive intermediates). Kedalo et al.[103] have performed MD simulations to show that the activation of highly excited N molecules on Ru surfaces during NH synthesis can be approximated by the Fridman–Macheret model.[104] The interatomic interactions were described by a ML potential, utilizing smooth overlap of the atomic positions (SOAP) descriptors,[105] to enable ab initio accuracy but with increased computational efficiency. Bal and Neyts[106] have demonstrated that the Fridman–Macheret model is insufficient to describe the vibrational nonequilibrium (i.e., lowering of the free energy barrier) during the catalytic dissociation of H and CH on Ni surfaces at low temperatures of the order of 750 K. The multitudes of potential reaction pathways in between two metastable states were simplified by applying a harmonic linear discriminant analysis as a supervised dimensionality reduction algorithm.[107] The states themselves were studied by conducting short MD simulations with enhanced sampling methods (i.e., variationally enhanced sampling[108], metadynamics[109, 110]) to introduce the vibrational nonequilibrium.
@@
Siron et al.[111] have proposed a high-throughput workflow for the study of amorphous material (i.e., a-C, a-Si, a-SiO, a-AlO) surface reactions relevant to dry plasma etching. Surface sites were characterized with SOAP descriptors and clustered with a Bayesian Gaussian Mixture model to effectively reduce sites to be studied for reaction kinetics, specifically etching.
@@
The bombardment of surfaces with energetic ions spawns collision cascades in the subsurface region, which eventually lead to persistent radiation damage and the emission of surface material atoms from the surface. The latter is referred to as sputtering and is of major importance for the sputter deposition of thin films as well as etching of high aspect ratio features.
@@
Hamedani et al.[112] have used an ML interaction potential to conduct an MD study on the radiation damage and sputtering of Si. The utilized ML potential, specifically the Gaussian approximation potential (GAP),[113] was originally trained to equilibrium properties and, hence, complemented with a DFT repulsive potential (i.e., DMol[114]) for short range interactions to resolve the evolution of the collision cascade accurately. The simulations provided sputtering yields that were in good agreement with experimental reference values. A greater number of isolated small defect clusters was observed than provided by MD simulations based on traditional interaction potentials (i.e., Stillinger-Weber[115], Tersoff III[116]). Moreover, a new defect structure was revealed, that is three split interstitials surrounding a central vacancy. This structure and the likelihood for its formation was further analyzed in a follow-up study.[117]
@@
Sputtering yields can also be predicted by utilizing the Yamamura model.[118] Phadke et al.[119] have showed that its model parameter Q, which relates to linear scaling, has a simple positive correlation with the sputter threshold energy. This result was obtained by applying a Bayesian Markov Chain Monte Carlo algorithm to TRIDYN[120] simulations of Ar and Ne ions bombarding 70 elemental targets. Eckstein and Preuss[121] applied Bayesian probability theory to improve the accuracy of the Yamamura model’s sputtering yield predictions by empirical means. Preuss et al.[122] conducted SDTrimSP[120, 123] simulations of D ions sputtering Fe targets to revise the model’s internal parameter (i.e., surface binding energy). This was achieved by to applying a reduced-order spectral expansion[124] uncertainty quantification based on the Bayesian framework.